OpenStudy (anonymous):

I need to make a table of values for 2x+9 if x>-2.

OpenStudy (mathstudent55):

For each value of x you see in the table, calculate 2x + 9 and place its value in the appropriate place in the table. 

OpenStudy (mathstudent55):

For x = -1, 2x + 9 = 2(-1) + 9 = -2 + 9 = 7 Now do the rest for the other values

OpenStudy (anonymous):